OpenStudy (anonymous):

Evaluate log 6. Round your answer to five decimal places.

OpenStudy (anonymous):

round ans 5 decimal places

OpenStudy (whpalmer4):

We're assuming the logarithm base 10, correct? Is there any more information given?

OpenStudy (doc.brown):

\(log_6(\)anything here?\()=\) or here?

OpenStudy (anonymous):

@whpalmer4 Evaluate common and natural logarithms using your calculator. Round your answer to five decimal places. log 4

OpenStudy (anonymous):

this is how its written

OpenStudy (anonymous):

@doc.brown no

OpenStudy (jdoe0001):

\(\Large{ \bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ \quad \\ log(6)\implies log_{{\color{red}{ 10}}}({\color{blue}{ 6}})\implies ?}\)

OpenStudy (jdoe0001):

hmmm actually that may not give us much anyway so... I gather you'd just have to use the calculator =)

OpenStudy (anonymous):

all it gives me its 0.77815125

OpenStudy (jdoe0001):

that's correct

OpenStudy (anonymous):

after that what do i do

OpenStudy (jdoe0001):

be happy =)

OpenStudy (anonymous):

i have to round ans 5 decimal places how do i do that

OpenStudy (anonymous):

I am happy :D

OpenStudy (jdoe0001):

ohh ahemm then round to 5 decimal places :) \(\bf 0.77815\quad 125\qquad \textit{5 decimal places}\)

OpenStudy (anonymous):

is it 0.78?

OpenStudy (jdoe0001):

well, you need 5 decimal places 0.78 <---- 2 places ^^ 0.77815125 <--- 5 places ^^^^^

OpenStudy (whpalmer4):

If you happen to know the values of some other logs, you could work it out without a calculator. For example, \[\log 6 = \log 2*3 = \log 2 + \log 3 \approx 0.30103 + 0.47712 = 0.77815\] If you do a lot of calculation, you'll find it convenient to memorize logs, square roots, etc. of a handful of values.